How much time is required to lift a 525 kg object 25.0 m with a motor that produces 2.50 kW?

Answer by scott8148(6628)

the force required to lift the object is equal to mg ___ f = 525 * 9.8

the work done is the force acting over distance ___ w = 525 * 9.8 * 25

power is work per unit of time ___ (525 * 9.8 * 25) / t = 2500

t = (525 * 9.8 * 25) / 2500

the units are all in the MKS system , so "t" is in seconds